Sol Summary – April 24th

Crew 297 Sol Summary Report 24-04-2024

Sol: 10

Summary Title: The Martian Chronicles

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning a portion of the crew woke up early to evade the impending hot weather and venture out on an EVA to Somerville Overlook. With the morning breeze assisting, the crew took the brisk walk from Opportunity to the overlook and admired the views. But of course, research was the main priority, and as such the crew went to work. Many samples and measurements were collected from the area, and on the way back Case was finally able to return to areas outside of the MDRS station.

Meanwhile, the later EVA crew, within the warming weather, was able to finally find the hidden NPS and return it safely to the Hab. The crew of two then set foot to Skyline Rim and investigated NPS digging sites along the way, testing the Martian surfaces for a qualitative assessment of ease of digging. This particular EVA broke a record for longest walking distance during Mission 297 – a Brobdingnagian 10 km!

Look Ahead Plan:

Tomorrow will be the last day involving multiple EVAs. The morning EVA will venture out to the Gateway of Lith and take Case out for its last stroll into Martian regolith. Case is expected to now be fully functional and should reliably take imagery of the area (with some expected intervention via crew manual driving). The second EVA will venture out to Somerville Overlook to scope areas for NPS digging. Back at the MDRS station, the crew will prepare the decommissioning process of the GreenHab for the season.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal. Inop’d for the current mission.
Weather:
Morning (data from 0900): calm winds (3 mph), chilly (59F), and sunny with clear skies
Noon: moderate winds (12 mph), warm (76F), and sunny with clear skies
Afternoon (1800): moderate winds (14 mph), hot (81F), and sunny with clear skies
Crew Physical Status: Overall crew health is nominal.

EVA: There were two EVAs today. The first EVA, involving Sarah and Pawel, successfully reached the Somerville Overlook and took gamma-ray spectrometer measurements along with samples. On the way back, near Compass Rock, the crew performed the human-interaction study with Case, though had to cut short due to time constraints. The second EVA was able to (finally) successfully locate the NPS and bring it back to the Hab safely. The crew also ventured to Skyline Rim and located many potential NPS digging locations on the way.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (2 EVAs)
EVALink Feedback Report
Support Requested: None requested.

Sol Summary – April 25th

Crew 297 Sol Summary Report 25-04-2024

Sol: 11

Summary Title: Earth Rising on the Red Horizon

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning, the first of two EVAs was conducted. Drs. Storch and Sawicki ventured out far north to the Gateway to Lith this morning to locate an area suitable for Case to drive around. Although they did not find anything in the vicinity of Gateway to Lith for this purpose, they admired the views of this area. On the way back, an adequate area was indeed discovered and utilized for the human-research interaction study. Case performed well in its given task, requiring minimal human intervention. The robot was then driven manually to test its steep hill-traversing abilities. The equipment was gathered up and the crew returned home to the Hab.

The second, late afternoon EVA, involved the two Crew Engineers Storch and Laude.This EVA team journeyed to Somerville Overlook to acquire telephoto photos of the sedimentary rock layers in the walls of the escarpments. As the weather started to head south, the team headed back towards MDRS, but made a planned stop at Chandor Chasma for additional photos of stratigraphy and rocks. Terrain was also examined for NPS recovery and was found extremely difficult for such an operation.

For dinner the crew was treated to a flavorful vegetable stir fry prepared by Chef (and GHO) Marquez who used a substantial amount of harvested greens from the GreenHab.

Look Ahead Plan:

Tomorrow is unfortunately our last sol at MDRS. We will start the day with a geological expedition EVA. When we return we will have our hands full with end-of-mission tasks: finishing reports, dismantling the GreenHab for the season, taking inventory, packing for our trip back to Earth and cleaning.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal. Inop’d for the current mission.
Weather:
Morning (data from 0900): calm winds (2 mph), chilly (59F), and scattered cloud coverage
Noon: calm winds (7 mph), warm (70F), and scattered cloud coverage
Afternoon (1800): moderate winds (16 mph), chilly (69F), and scattered cloud coverage
Crew Physical Status: Overall crew health is nominal.

EVA: There were two EVAs today. The first EVA had a successful testing effort of Case on the way back from the scenic Special Region and Gateway to Lith. The second EVA went to Somerville Overlook and Candor Chasma, collecting telephotos of the geological features in those areas for review by the crew geologist.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (1 EVA)
Support Requested: None requested.

Sol Summary – April 26th

Crew 297 Sol Summary Report 26-04-2024

Sol: 12

Summary Title: So Long, And Thanks for All the Fish

Author’s name: Pawel Sawicki

Mission Status: Concluded

Sol Activity Summary:

This morning the crew woke up with a heartfelt bittersweet feeling about the approaching trip back to Earth. All of the tasks and duties that we habitually had taken for granted started becoming "lasts." One of those lasts was the final EVA. Lamm, Lynch, Storch, and Sawicki participated in a geologoical expedition of North Ridge. With a great view of our Martian home, we reminisced the past 12 sols and investigated the surrounding kaleidoscope of varying rock colors. Fortunately for the EVA crew, it began to rain just as the crew returned to the airlock.

At the Hab, we had our hands full with end-of-mission tasks: finishing reports, dismantling the GreenHab for the season, taking inventory, packing for our trip back to Earth, and cleaning. Our last dinner was prepared by Chef Lynch, a wonderfully crafted pair of pot-pies. The crew will spend their last night together just as they have done for the other 12: enjoying each others’ company.

Look Ahead Plan: Returning to our home planet.

Anomalies:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal.
Suit 9 (broken visor)
Weather:
Morning (data from 0900): calm winds (0 mph), chilly (55F), and overcast
Noon: moderate winds (17 mph), chilly (60F), and overcast
Afternoon (1800): moderate winds (14 mph), chilly (61F), and overcast with intermittent rain
Crew Physical Status: Overall crew health is nominal.

EVA: There was one EVA today. The crew of four traversed up North Ridge, where green shale, tan sandstone, and various colors of regolith were examined. This EVA marked the 18th and final EVA of Mission 297.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
Mission Summary Report
End-of-mission Research Report
Food Inventory
Support Requested: Stop by the Hab before we leave at 11 am tomorrow, so we can thank you in person for all your help!

Sol Summary – April 23rd

Crew 297 Sol Summary Report 23-04-2024

Sol: 09

Summary Title: Just Keep Searching

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning, three crew members set foot outside of the Hab with one goal in mind: finding the NPS. They scurried across the rugged terrain with their advanced searching technologies and methodologies, diligent eyes, and unwavering dedication to locate a simple, yet elusive object. They searched and searched, giving it all they had. Yet, all of their efforts were futile. The NPS remained unfound and waiting, optimistically perhaps, to be found on another EVA on another day.

Fortunately, the crew was not left empty handed the entire day. Case the robot displayed enough capabilities today to let him out on another field deployment. With on-lookers both inside and outside the ScienceDome, Case was able to relatively smoothly (with some manual intervention) traverse an obstacle course laid out before it. Once the crew gained confidence in its abilities, they used Case to recreate a part of "The Martian": generating a hexadecimal polar grid to relay messages along.

All crew members, regardless of how successful their EVAs turned out, were treated to flavorful Southwest-style beans, rice, and veggies topped with Fritos, as prepared by Chef Laude.

Look Ahead Plan:

Tomorrow will be composed of two EVAs. The morning EVA will involve a geological expedition to the orange (via satellite imagery) regions. This EVA will also have additional testing with Case with the objective of capturing valuable 360 degree imagery of any significant geological sites. The second EVA team will do their best to evade the morning sun and find the unfound NPS from today, while scoping out suitable regions for NPS excavation research.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal. In-work.
Weather:
Morning (data from 0900): calm winds (0 mph), chilly (59F), and sunny with clear skies
Noon: calm winds (3 mph), warm (76F), and sunny with clear skies
Afternoon (1800): strong winds (20 mph), hot (83F), and scattered cloud coverage
Crew Physical Status: Overall crew health is nominal.

EVA: There were two EVAs today. The first EVA was unsuccessful in finding the NPS after an arduous search effort. Precise coordinates have been given of the NPS and a consequent recovery EVA is planned tomorrow. The second EVA scrutinized the capabilities of Case further, and while there were setbacks, Case has been deemed ready for field deployment.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (2 EVAs)
Support Requested: We request that the 1-hour power blackout period tomorrow be from 1600-1700. We also request support to pick-up (Qty. 2) bags of No-Burn Waste. Thanks in advance for your help.

Sol Summary – April 20th

Crew 297 Sol Summary Report 19-04-2024

Sol: 06

Summary Title: Live Another Sol

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Our Sol 6 went quite a bit smoother than the infamous Sol 6 of Ares III and Mark Watney.

The crew started the sol feeling rejuvenated after some great sleep, induced by indulging on the delicious late-night apple pie made by Matt Lynch. Shortly after waking up in the morning, the crew started getting ready for their 4-hour EVA. They were provided coordinates of the desired landing spot for the NPS, though strong winds last night increased the NPS search radius to 400 meters. Fortunately for the crew, it took only several minutes to find the NPS and carry it back to the rovers. This left more than three hours for the crew to pursue objectives for their respective research projects. Sarah utilized the additional time to log gamma-ray spectral data of samples within the beautiful landscape. The Case robot unfortunately was quite shy today, and did not want to communicate with the rest of the crew, consequently shelving his objective for another day. "I guess you could call it a failure," but I prefer the term "learning experience" (Watney, in The Martian, Weir). With an additional crewmate available for excavating, a hole adequate for the NPS was dug in 67 minutes, even through a challenging surface.

The crew came back from their EVA to freshly-made scones, prepared by Dave. The rest of the afternoon was spent performing nominal duties, conducting and preparing research, and writing reports. Dinner was made by Chef Marquez, a refined shoyu ramen recipe exquisitely composed of a dashi-base crafted with kombu seaweed, shiitake mushroom, and tare. The ramen was then brought together with an egg-based ramen, kanpyo, flavored tomato, onion, and GreenHab-harvested red cabbage and green onion. The umami-packed, harmonious mixture came together to bring us all closer together in the later hours of the evening.

Look Ahead Plan:

Tomorrow will be a rest day of sorts for the crew with a late wake-up and less work planned than usual. Nevertheless, there are two low-activity EVAs planned tomorrow. The first EVA will involve Team Matt picking a suitable location for the NPS and hiding it for a second team to recover. The second EVA will be more focused around the station, specifically at the RAM. Given Case’s difficulties today, the robot will be tested right outside of the Hab. Dave will manually drive the robot from a ground station in the RAM while Pawel monitors closeby.

Anomalies in work:

Suit 11 (fan power significantly decreased mid-EVA): In work
GreenHab automated fan: Work-around has been implemented to power on the fan manually. In case of power outage or inoperable fan, the GHO will open the door for cooling.
ScienceDome: We have noticed a foul chemical smell the last few days which has made the ScienceDome unsuitable for any work.
EVA headset #2 not receiving comms when connected with two way radio: R&R with spare headset.
Weather:
Morning (data from 0900): calm winds (2 mph), chilly (57F), and sunny with clear skies
Noon: calm winds (2 mph), warm (68F), and few clouds in the sky
Afternoon (1800): moderate winds (19 mph), warm (74F), and sunny with clear skies
Crew Physical Status: EVA Crew members are physically tired from the EVA today, but recovering well. Overall crew health is nominal.

EVA: Today’s EVA was nominal and an overall success. The crew of 4 (Sawicki, Storch, Lynch, Lamm) was able to locate the hidden NPS quickly today, within 10 minutes of parking the rovers and taking off on foot. The crew continued with their secondary mission objectives. Gamma-ray spectral data was collected from samples in the area and the crew was able to dig (and fill) a hole suitable for the NPS. Unfortunately, Case had network issues and was not utilized today.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (2 EVAs)
Mid-Mission Report
Support Requested: None requested.

Sol Summary – April 21st

Crew 297 Sol Summary Report 21-04-2024

Sol: 07

Summary Title: I’m Sorry Dave, I’m Afraid I Can’t Do That

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

The crew woke up today well rested after a night of watching the Martian and getting to sleep in, a celebration for passing the mid-mission mark. Today’s docket of EVAs consisted of i) the Matts adventuring far northwest searching for an NPS hiding spot around the Sea of Shells and ii) Case adventuring right outside of the RAM. Both EVA crews were satisfied with what they saw today, whether it was the spectacular views of the Overlook or a fully functioning User Interface. The whole crew was later treated to a bon gou Jambalaya by Chef Lamm, a dish reminiscent of New Orleans street jazz – soulful, vibrant, and complex characteristics heard within each bite.

Look Ahead Plan:

Tomorrow there will only be one EVA, commencing shortly after the Martian sun wakes up. During this EVA, Sarah and Pawel intend on exploring interesting geological features to the northwest of the Hab within the vicinity of the Sea of Shells and Overlook. The rest of the crew will rest from EVAs and continue their nominal duties including preparing and conducting IVA research.

Anomalies in work:

Suit 11 (fan power significantly decreased mid-EVA): Battery has been replaced.
GreenHab (automated fan not working): working now that generator power is back online
ScienceDome (foul smell): Source of odor (batteries) has been determined and removed. Crew will continue to monitor for another day before using the ScienceDome fully.
Weather:
Morning (data from 0900): calm winds (2 mph), chilly (57F), and sunny with clear skies
Noon: calm winds (2 mph), warm (68F), and sunny with clear skies
Afternoon (1800): moderate winds (14 mph), warm (77F), and sunny with clear skies
Crew Physical Status: Crew health is nominal.

EVA: There were two EVAs today with two crewmates each – a first for us. The first EVA crew set out to hide the NPS, but due to a lack of good hiding spots in the area this goal was delayed to a later EVA. The crew remarked on the spectacular landscapes that part of Mars has to offer. The other EVA was more local to the Hab (by the RAM), where Case went out for a walk to test its features. Operation of the robot was as smooth as expected and more challenging EVAs are to come for Case.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (1 EVA)

Support Requested: None requested.

Sol Summary – April 22nd

Crew 297 Sol Summary Report 22-04-2024

Sol: 08

Summary Title: Mars on Earth Day

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning, half the crew woke up with the Martian sun to prepare for an early EVA start time. Keeping cool prior to the impending peak of the hottest sol of the sim, Lamm and Sawicki set out to survey the grey area (as seen by satellite) between the Sea of Shells and Overlook to gather spectral data. Unfortunately for them, Opportunity was only able to traverse half of Brahe Highway and the rest of the journey was conquered by foot. While the views of Overlook proved to be spectacular, it was also confirmed that the area contained more uranium content than the previously explored lands around MDRS (but still in line with an average amount). As the environment started to heat up, the crew forwent a visit to the Sea of Shells and headed back to the Hab. After conducting their duties today, the crew enjoyed a finger-licking tuna & veggie casserole made by Chef Lynch.

Look Ahead Plan:

Tomorrow there are a couple of EVAs planned. The early EVA crew will look to locate the NPS that was hidden today. The second EVA will continue locally testing Case, with a ground station located inside the comfortable air-conditioned confines of the ScienceDome. During the latter EVA, the crew will monitor Case on the outside and take gamma-ray spectral data of geological formations by the MDRS station.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal
Weather:
Morning (data from 0900): calm winds (2 mph), chilly (57F), and sunny with clear skies
Noon: calm winds (4 mph), warm (75F), and sunny with clear skies
Afternoon (1800): moderate winds (19 mph), warm (80F), and broken cloud coverage
Crew Physical Status: Overall crew health is nominal.

EVA: During today’s EVA the crew was able to admire the beauty of the Overlook, but unfortunately the crew was not able to reach the Sea of Shells. Nevertheless, site-seeing is never an objective of any EVA, including today’s. Thus the goals of the EVA (analyzing the geological features of this Northwest area) were completed and the EVA can be deemed a success.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (2 EVAs)
Mid-Mission Report
Support Requested: None requested.

Sol Summary – April 17th

Crew 297 Sol Summary Report 17-04-2024

Sol: 03

Summary Title: Red Planet Rocks!

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary: The crew started the day at 0800, welcomed by the smell of the delectable crepes prepared by Sean and Matt S. With the mood already in good spirits after breakfast, the EVA team could not flounder, as they hastily were able to locate the NPS and spent the rest of the time analyzing the rocks of the Red Planet. After a fortunate EVA, the crew was treated to a mouthwatering take on fromage et legumes soup prepared by Matt L. While the soup was prepared, the crew learned how to play the cooperative card game Hanabi. In the afternoon, Sean and Pawel troubleshooted thermal problems within the GreenHab along with harvesting a ripe cherry tomato. Dinner concluded a day of scrumptious meals, with a soy peanut couscous containing vegetables and chicken, prepared by chefs Matt L. and Sean.

Look Ahead Plan:

Tomorrow we will perform a single EVA in the morning. This EVA will be the first for Case, our robotic crewmate, who will traverse and record imagery as we monitor closely. The EVA will also test out digging possibilities for the [inert] NPS.

Anomalies in work:

Suit 5 (ventilation electrical connection): Investigated and likely fixed. To be further examined/tested in the coming days.
GreenHab automated fan: The GreenHab fan powered on today when the GH temp reached 94F, but then turned off. It then powered cycled continuously but never reached its expected rpm.
Suit 4 (collar cable ring missing): In work
Weather:
Morning (data from 0900): calm winds (2 mph), chilly (53F), and clear skies (0% cloud coverage)
Noon: moderate winds (15 mph), warm (70F), and clear skies (0% cloud coverage)
Afternoon (1800): moderate winds (17 mph), warm (76F), and clear skies (0% cloud coverage)
Crew Physical Status: Crew health is nominal.

EVA: EVA was nominal today. The crew found the hidden NPS fairly quickly (~15 minutes) and spent the rest of the time performing gamma-ray spectrometer readings of regions along Cow Dung Road south of MDRS including Robert’s Rock Garden.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (1 EVA)
Support Requested: We request that the [inert] NPS be placed by Mission Support at a location of their choosing (though reachable by rover and within a maximum of 15 minutes of walking) in either Candor Chasma, Eos Chasma, Valles Marineris, or Aurorae Chaos for a search and recovery EVA on 19-04-2024. The NPS will be placed in the rear airlock tomorrow night, 18-04-2024 by 1600. Once the inert NPS is hidden, precise GPS coordinates and a photo are requested to be taken to not lose track of the location. These GPS coordinates are then asked to be sent to Pawel.

Sol Summary – April 18th

Crew 297 Sol Summary Report 18-04-2024

Sol: 04

Summary Title: stanleY yelnatS tuanortsA

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Today we opened our eyes to a welcomed Martian morning partially filled with clouds. These clouds proved to be fortuitous during the later EVAs in reducing sun glare on screens and the need for respiration during digging. Today’s EVA was a first for our robotic crewmate Case. Case, roughly the size of its distant Martian cousin, Sojourner, has the intention of helping us take useful Martian imagery for planning purposes of future crews. Today was also the first time we had four crewmates involved in the same EVA and the first true attempt at digging the NPS. Case was able to capture useful imagery of the surroundings, yet, it was not able to take off its metaphorical leash, needing to be monitored constantly and manually driven for a large portion of its adventure. The other two EVA crew mates had great success – digging a hole suitable for the NPS at a rate of 30 cubic feet per hour. Two hours had passed before the rover and related GSE were collected, the hole was filled up, and we started heading back home. When the crew finally recovered from the eventful EVAs they learned about the functionality of WIDGIT and how to incorporate it into their collaborative report writing workflows. FInally, a savory Japanese curry dish was prepared and served by Chef Sean to the delight of the crews’ taste buds.

Look Ahead Plan:

Tomorrow we will have a chock-filled day of EVAs from 0900 to 1600. The first EVA will look to locate the NPS that has thankfully been placed by Mission Support for us. After they successfully find the NPS, they will provide the favor to the second EVA crew of hiding the NPS. The second EVA crew will then attempt to find the NPS with enough time to spare for additional gamma-ray spectral readings of the Martian rocks in the surrounding area.

Anomalies in work:

Suit 5 (ventilation electrical connection): Investigated and likely fixed. To be further examined/tested in the coming days.
GreenHab automated fan: The GreenHab fan powers on when the GH temp reaches 94F, but then shortly thereafter turns off. It power cycles continuously but never reaches its expected rpm.
Suit 4 (collar cable ring missing): Fixed today, will be functionally tested during EVA tomorrow.
ScienceDome: We have noticed a foul chemical smell. Dispositioning.
Weather:
Morning (data from 0900): calm winds (6 mph), chilly (57F), and partially cloudy
Noon: calm winds (3 mph), warm (67F), and sunny
Afternoon (1800): calm winds (8 mph), warm (74F), and broken cloud coverage
Crew Physical Status: Crew health is nominal.

EVA: Nominal EVA. The crew of four involved with the EVA was able to test digging capabilities within their spacesuits and use the rover, Case, to capture MDRS site imagery.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (1 EVA)
Support Requested: We have requested that the [inert] NPS be placed by Mission Support at a location of their choosing (though reachable by rover and within a maximum of 15 minutes of walking) in either Candor Chasma, Eos Chasma, Valles Marineris, or Aurorae Chaos for a search and recovery EVA #6 on 19-04-2024. The NPS will be placed in the rear airlock tomorrow night, 18-04-2024 by 1600. Once the inert NPS is hidden, precise GPS coordinates and a photo are requested to be taken to not lose track of the location. These GPS coordinates are then asked to be sent to Pawel.

We also request to have the permission for Opportunity to be plugged in after EVA #6 tomorrow for one hour (1200-1300) between EVAs to extend its range.

Sol Summary – April 19th

Crew 297 Sol Summary Report 19-04-2024

Sol: 05

Summary Title: So much universe, and so little time

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Today we all woke up early to the sounds of Strauss and the smells of freshly made pancakes by Chef Laude. Shortly thereafter the first of two EVAs commenced, where, within an hour of leaving the airlock, Matt S. located the [inert] Nuclear Power System (NPS). After the crew hid the NPS they met all their EVA objectives and returned to the Hab. The second EVA group was not quite as swift as the first in finding the NPS, but nonetheless Pawel was able to locate the NPS within an hour of EVA start. Collecting gamma-ray spectral data was unfortunately cut short during the second EVA due to a more urgent matter: a nonconforming fan on Suit 11, resulting in an expedited return back to the Hab. The day concluded with a spaghetti dinner prepared by Chefs Laude and Storch, report writing, and late-night games.

Look Ahead Plan:

Tomorrow we have an eventful EVA planned with three of the crew’s research topics consolidating into one EVA objective. The crew of four (Sawicki, Storch, Lynch, and Lamm) will first locate the [inert] NPS. Once found, the crew will use Case (which has recently been equipped with a Geiger counter) to approach the [inert] NPS and verify that the unit is safe to approach. During the recovery efforts, continued use of the gamma-ray spectrometer will further characterize the Martian rocks.

Anomalies in work:

Suit 11 (fan power significantly decreased mid-EVA): EVA was abruptly cut short due to low fan rpm, causing pressing concerns for heat exhaustion. Prior to the 1.5 hour EVA, the suit battery read 13.5 V and afterwards read 11.9 V.
GreenHab automated fan: The GreenHab fan powers on when the GH temp reaches 94F, but then shortly thereafter turns off. It power cycles continuously but never reaches its expected rpm.
ScienceDome: We have noticed a foul chemical smell the last few days which has made the ScienceDome unsuitable for any work. This intense odor has assuaged during the current sol. We still continue to monitor closely.
EVA headset #2 not receiving comms when connected with two way radio (radio works fine).
Weather:
Morning (data from 0900): calm winds (1 mph), chilly (59F), and few cloud coverage
Noon: calm winds (2 mph), warm (69F), and scattered cloud coverage
Afternoon (1800): moderate winds (17 mph), warm (71F), and broken cloud coverage
Crew Physical Status: Sarah is feeling a bit fatigued, but recovering well. The rest of the crew is presently nominal in terms of health.

EVA: The first EVA today was a complete success. The crew was able to locate the NPS and re-hide it for the next crew in great time. While the second EVA crew was able to locate the NPS quite quickly (~25 minutes), there was a fan malfunction within suit #11 that resulted in an aborted EVA and expedited return to the Hab. Sarah had begun exhibiting symptoms of heat exhaustion from the lack of airflow within her helmet.

To loosely paraphrase Artemis by Weir: "[Mars] is mean. [He] doesn’t care why your suit fails. [He] just [cancels your EVA] when it does."`

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (1 EVA)
Support Requested: We have requested that the [inert] NPS be placed by Mission Support at a location of their choosing (though reachable by rover on Galileo Road 1104 and within a maximum of 25 minutes of walking) for a search and recovery EVA #8 on 20-04-2024. The NPS will be placed in the rear airlock tonight, 19-04-2024 by 2000. Once the inert NPS is hidden, precise GPS coordinates and a photo are requested to be taken to not lose track of the location. These GPS coordinates are then asked to be sent to Sean. Many thanks for the support provided thus far for the NPS placements!

We have also requested one additional EVA headset, as one is not functioning properly.

Copyright © The Mars Society. All rights reserved. | Main Site